how do you figure out the equation for rectangular numbers?
ex. the first term is 1, there are 2 dots, the second term is 2, there are 6 dots, the third term is 3, there are 12 dots.
I see that the first difference is 2 more than the previous, and the second difference is 2, but how do i write this in an equation? Thank you.

Answer by stanbon(75887)

n(n+1) gives the sequence
1*2=2
2*3=6
3*4=12
4*5=20
etc.
=========
Cheers,
Stan H.